Anthracene is a yellow, crystalline solid found in coal tar. Complete this structure for anthracene, C14H10, by adding bonds and hydrogen atoms as necessary.


Anthracene is a yellow, crystalline solid found in
Answer #1
Concepts and reason

Draw the structure of anthracene and count the number sigma bonds, number of pi bonds, and number of valence electrons. Also identify the hybridization of carbon atoms present in anthracene.

Fundamentals

Identify the hybridization of an atom by counting the number of sigma bonds.

If carbon is attached with 4 sigma bonds then the hybridization is sp3{\rm{s}}{{\rm{p}}^3} .

If carbon is attached with 3 sigma bonds then the hybridization is sp2{\rm{s}}{{\rm{p}}^{\rm{2}}} .

If carbon is attached with 2 sigma bonds then the hybridization is sp{\rm{sp}} .

The molecular formula of anthracene is C14H10{{\rm{C}}_{{\rm{14}}}}{{\rm{H}}_{{\rm{10}}}} .

Calculatethedoublebondequivalenceasfollows:DBE=NC+1((NH+NCl)NN2)ThemolecularformulaforthegivencompoundisC14H10DBE=14+1((10+0)02)DBE=10So,thecompoundhas10doublebonds,itindicatesthepresenceofbenzeneringintheanthracene.\begin{array}{l}\\{\rm{Calculate}}\,\,{\rm{the}}\,\,{\rm{double}}\,\,{\rm{bond}}\,\,{\rm{equivalence}}\,\,{\rm{as}}\,\,{\rm{follows:}}\\\\{\rm{DBE}}\,{\rm{ = }}\,{{\rm{N}}_{\rm{C}}}\,{\rm{ + }}\,{\rm{1}}\, - \left( {\frac{{{\rm{(}}{{\rm{N}}_{\rm{H}}}\,{\rm{ + }}\,{{\rm{N}}_{{\rm{Cl}}}}{\rm{)}} - {{\rm{N}}_{\rm{N}}}}}{{\rm{2}}}} \right)\\\\{\rm{The}}\,{\rm{molecular}}\,{\rm{formula}}\,{\rm{for}}\,{\rm{the}}\,{\rm{given}}\,{\rm{compound}}\,{\rm{is}}\,\,{{\rm{C}}_{14}}{{\rm{H}}_{{\rm{10}}}}\\\\{\rm{DBE}}\,{\rm{ = }}\,14{\rm{ + 1}}\, - \left( {\frac{{{\rm{(10}}\,{\rm{ + }}\,{\rm{0)}} - {\rm{0}}}}{{\rm{2}}}} \right)\\\\{\rm{DBE}}\,{\rm{ = }}\,\,10\\\\{\rm{So,}}\,{\rm{the}}\,{\rm{compound}}\,{\rm{has}}\,\,10\,{\rm{double}}\,{\rm{bonds,}}\,{\rm{it}}\,{\rm{indicates}}\,{\rm{the}}\,{\rm{presence}}\,{\rm{of}}\,{\rm{benzene}}\,{\rm{ring}}\\\\{\rm{in}}\;{\rm{the}}\;{\rm{anthracene}}.\\\end{array}

The structure of anthracene is,

н
на
с
с
Н
н
н

Hybridization is sp2{\rm{s}}{{\rm{p}}^{\rm{2}}}

Number of sigma bonds is 26.

Number of pi bonds is 7.

Number of valence electrons in sigma bonds is 52 electrons.

Number of valence electrons in pi bonds is 14 electrons.

Ans:

sp2{\rm{s}}{{\rm{p}}^{\rm{2}}} hybrid orbitals are utilized by carbon in anthracene.

26 sigma bonds and 7 pi bonds are present in anthracene.

The number of valence electrons in sigma bond are 52 and number of pi electrons are 14.